OpenStudy (anonymous):

How can you find t from this equation? :

OpenStudy (anonymous):

\[(2)^{t/3} = (3)^{t/2} + 10,000\]

ganeshie8 (ganeshie8):

graph it maybe ?

OpenStudy (anonymous):

can't we use any of the logs or something? :)

ganeshie8 (ganeshie8):

logs are of no use here, you mean below right ? \(\large (2)^{t/3} = (3)^{t/\color{red}{5}} + 10,000 \)

OpenStudy (anonymous):

wait, haha, my brain is loading.. why's it 5? :) red

ganeshie8 (ganeshie8):

population A doubles in 3 hours population B triples in 5 hours you're stuck wid that problem eh ?

OpenStudy (anonymous):

no, in 2 hours

OpenStudy (anonymous):

no, i already got the answer for that problem. it's not 10,000 more. it's 10,000 times more.

ganeshie8 (ganeshie8):

oh ! then its real easy

ganeshie8 (ganeshie8):

\(\large (2)^{t/3} = (3)^{t/2} + 10,000 \)

OpenStudy (anonymous):

typographical error at first. so the equation for my first problem was this: \[(2)^{t/3} = (3)^{t/2} * 10,000\] and i got the answer from using natural log according to one person who replied

OpenStudy (anonymous):

but im just curious if the problems says, 10,000 more and not 10,000 times more.

ganeshie8 (ganeshie8):

yeah u can log and work it if its 10000 times

ganeshie8 (ganeshie8):

good, think about it :) let me knw if u find something interesting =))

ganeshie8 (ganeshie8):

hey wait a sec, \(\large (2)^{t/3} = (3)^{t/2} * 10,000\)

OpenStudy (anonymous):

ahmmm, *brainstorming* haha. brainstorming with an empty brain. my brain is loading right now lol okay i'll tryyy

ganeshie8 (ganeshie8):

this is impossible right ? left side is growing very slow. so it can never reach right side

ganeshie8 (ganeshie8):

empty brain is always good, cuz u can fill it wid nice stuff... :)

OpenStudy (anonymous):

really? it's impossible? so it's my teacher's fault why students can't get it. waoh

ganeshie8 (ganeshie8):

population A doubles in 3 hours population B triples in 2 hours B is way faster than A. so A cannot catchup B

OpenStudy (anonymous):

maybe someday it will LOL

ganeshie8 (ganeshie8):

i remember, in ur other question, you have different numbers :- population A doubles in 3 hours population B triples in \(\color{red}{5}\) hours this is possible however.

OpenStudy (anonymous):

wait, goin' to check that

ganeshie8 (ganeshie8):

okie :o

OpenStudy (anonymous):

@ganeshie8 typographical error again! hahahaha im so sorry, it's 5 =)

OpenStudy (anonymous):

so sorry so sorry so sorry hahaha

ganeshie8 (ganeshie8):

np :) just making sure.... you dont cuss ur teacher when there is no fault from her lol

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

so it's \[(2)^{t/3} = (3)^{t/5} +10,000\]

ganeshie8 (ganeshie8):